OpenStudy (anonymous):

A country's population in 1995 was 228 million. In 2001 it was 230 million. Estimate the population in 2010 using the exponential growth formula. Round your answer to the nearest million

OpenStudy (cwrw238):

P = P0 * e^(kt) here P = 230 m , P0 = 228 m , t = 2001 - 1995 = 6 from this you can find the constant k then you have the formula to find P in 2010

OpenStudy (cwrw238):

can you continue with this?

OpenStudy (anonymous):

1.008=e^k6 ln(1.008)=6k k=ln(1.008)/6

OpenStudy (anonymous):

so far right?

OpenStudy (cwrw238):

yes

OpenStudy (cwrw238):

maybe its better to use 1.0088 - not sure how much difference it would make though

OpenStudy (anonymous):

is it 231 million?

OpenStudy (cwrw238):

that would make k = 0.00146

OpenStudy (anonymous):

228e^.0135 ? 231 doesnt seem right.

OpenStudy (cwrw238):

so P at 2010 = 228 * e^(0.00146*15) = 233 million according to my calculation - ill re try

OpenStudy (anonymous):

k, yours seems more right.

OpenStudy (cwrw238):

yes i got same value again

OpenStudy (anonymous):

\o/ correct!